# Coreset Strikes Back: Improved Parameterized Approximation Schemes for (Constrained) $k$-Median/Means

### v2

@@ 4.2 $(\alpha,\beta)$-Fair Clustering and Variants
We begin with the definition of $(\alpha,\beta)$-fair $k$-median problem that has been studied in [CKLV17, BCJ+22, BFS24, HLLW25].
@@
The following definitions pertaining to coresets for $(\alpha,\beta)$-fair clustering were introduced in [BFS24].
Definition 16 (Coloring Constraint and Universal Coresets, [BFS24]).
@@
Universal coresets of size $\Gamma\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ in general metrics were designed in [BFS24], and the extensions to $(k,z)$-clustering variant (which is defined by naturally modifying the definitions above by including the $z$-th power of distances) with improved size were given in [HLLW25]. For our purpose, the former guarantee on the size suffices. Further, [BFS24] also designed an algorithm that takes input a weighted point set $(Y,w)$, and a set $X=\left\{c_{1},c_{2},\ldots,c_{k}\right\}$ of $k$ centers, runs in time $(k\Gamma)^{O(k\Gamma)}\cdot n^{O(1)}$, and returns an an assignment $f:Y\times X\to\mathbb{N}$, such that the cost $\sum_{p\in Y}\sum_{i\in[k]}f(p,c_{i})\cdot{\sf d}(p,c_{i})$ is minimized across all assignments that respect the $\alpha,\beta$ fairness constraints as above. Further, it is easy to check that this algorithm does not depend on the objective function – in fact it can compute an optimal assignment for any $(k,z)$-clustering objective. Thus, we have all the ingredients to plug into our framework for $(\alpha,\beta)$-fair $(k,z)$-clustering in metric spaces of bounded scatter dimension for any fixed $z\geq 1$, and obtain an EPAS with running time $\Gamma^{O(\Gamma k\lambda(\varepsilon)\log(\frac{1}{\varepsilon})(\log k+\log(\lambda(\varepsilon))))}\cdot(|\mathcal{I}|)^{O(1)}$.
@@
[BFS24] discuss the following variants of constrained clustering in their paper.
@@
[BFS24] discuss in detail the modifications (if any) to the universal coreset and the FPT algorithm for finding an optimal assignment required to handle each of these variants, and we direct the reader to these discussions. Since these are the only two ingredients required in our problem, our framework implies an EPAS for all of these variants in metric spaces of bounded scatter dimension.
4.3 Vanilla (Voronoi) Clustering
@@
First, let us suppose that there exists a point $p\in W$ such that ${\sf d}(p,x_{i})>\frac{6r^{\prime}_{i}}{\varepsilon}$ for all $i\in[k]$. Then, by arguments similar to that in 14.1, we can argue that ${\sf d}(p,X)<(1+\varepsilon){\sf d}(p,O)$, contradicting the assumption that $p\in W$. This shows part (i).
@@ 4.4 Additional Remarks
Jaiswal and Kumar [JK23] (also [DGI25]), extending the work of Agrawal et al. [AISX23] gave a general additive-$\varepsilon$ approximation-preserving reduction from constrained $(k,z)$-clustering with $m$ outliers to its analogous constrained version without outliers that has a multiplicative overhead of $\left(\frac{k+m}{\varepsilon}\right)^{O(m)}\cdot n^{O(1)}$. By plugging in our EPASes for constrained clustering into this result, we also obtain EPASes for the corresponding outlier versions.
@@
For most known metric spaces with bounded scatter dimension, the recent work of Huang et al. [HLLW25] gives coresets of size $\text{poly}(k,\varepsilon)$ for a large class of constrained $(k,z)$-clustering problems, including capacitated clustering, $(\alpha,\beta)$-fair clustering, fault-tolerant clustering, and more. To put this work into a proper context w.r.t. our EPASes, two remarks are in order.
Firstly, we note that such coresets can be used to obtain EPASes for the corresponding problems [^13] Although the following argument is very simple and should be folklore, we are unable to find a reference that explicitly describes how coresets of size independent of $\log n$ may be used to obtain EPASes for constrained $(k,z)$-clustering. Therefore, we provide the argument for the sake of completeness in order to perform a fair comparison with our results.. Let $(Y,w)$ be the coreset of size $\text{poly}(k,\varepsilon)$. Then, for each $p\in Y$ and each $i\in[k]$, we guess (up to a factor of $(1+\varepsilon)$) the portion of weight of $p$ assigned to the $i$th cluster. Note that the number of guesses is bounded by $(\frac{\log n}{\varepsilon})^{k\cdot|Y|}$, which is FPT in $k$ and $\varepsilon$. Fix one such guess. Then, for each cluster $i\in[k]$, we can find a feasible center that optimizes the cost (up to a factor of $(1+\varepsilon)$ arising from the guessing). Let $X$ be the set of centers chosen in this manner. We compute an optimal assignment using the assignment algorithm in time FPT in $k$ and $\varepsilon$. Finally, we return the minimum-cost solution found over all guesses. Although this strategy also yields an EPAS for many of the constrained problems in metric spaces of bounded scatter dimension, we note that the running time is $(k|Y|/\varepsilon)^{O(k|Y|)}\cdot n^{O(1)}$, and the size of coresets is $\Omega_{\varepsilon}(k^{2})$. Therefore, the dependence on $k$ of the running time is $2^{\Omega_{\varepsilon}(k^{3}\log k)}\cdot n^{O(1)}$, which is much worse compared to our running time of $2^{O_{\varepsilon}(k\log k)}\cdot n^{O(1)}$. Secondly, although the result of [HLLW25] gives coresets of size independent of $\log n$ in most known metric spaces of bounded scatter dimension, it is not known to yield such coresets for all metric spaces of bounded scatter dimension. Thus, this result does not answer the open question of [ABB+23] in its entirety.
@@
Acknowledgments.
We thank Andreas Emil Feldmann for bringing to our attention the work [FV22], which rules out the existence of an EPAS for the Capacitated $k$-Center problem in metrics of bounded highway dimension.

### v3

@@ 4.2.1 $(\alpha,\beta)$-Fair Clustering and Variants
We begin with the definition of $(\alpha,\beta)$-fair $k$-median problem that has been studied in [CKL+17, BCJ+22, BFS24a, HLL+25].
@@
The following definitions pertaining to coresets for $(\alpha,\beta)$-fair clustering were introduced in [BFS24a].
Definition 18 (Coloring Constraint and Universal Coresets, [BFS24a]).
@@
Universal coresets of size $\Gamma\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ in general metrics were designed in [BFS24a], and the extensions to $(k,z)$-clustering variant (which is defined by naturally modifying the definitions above by including the $z$-th power of distances) with improved size were given in [HLL+25]. For our purpose, the former guarantee on the size suffices. Further, [BFS24a] also designed an algorithm that takes input a weighted point set $(Y,w)$, and a set $X=\left\{c_{1},c_{2},\ldots,c_{k}\right\}$ of $k$ centers, runs in time $(k\Gamma)^{O(k\Gamma)}\cdot n^{O(1)}$, and returns an an assignment $f:Y\times X\to\mathbb{N}$, such that the cost $\sum_{p\in Y}\sum_{i\in[k]}f(p,c_{i})\cdot{\sf d}(p,c_{i})$ is minimized across all assignments that respect the $\alpha,\beta$ fairness constraints as above. Further, it is easy to check that this algorithm does not depend on the objective function – in fact it can compute an optimal assignment for any $(k,z)$-clustering objective. For GenBALL-INT, we use the underlying BALL-INT algorithm to compute a feasible center for every request set $Q_{i},i\in[k]$. Thus, we have all the ingredients to plug into our framework for $(\alpha,\beta)$-fair $(k,z)$-clustering in metric spaces of bounded scatter dimension for any fixed $z\geq 1$, and obtain an EPAS with running time $\Gamma^{O(\Gamma k\lambda(\varepsilon)\log(\frac{1}{\varepsilon})(\log k+\log(\lambda(\varepsilon))))}\cdot(|\mathcal{I}|)^{O(1)}$.
@@
[BFS24a] discuss the following variants of constrained clustering in their paper.
@@
[BFS24a] discuss in detail the modifications (if any) to the universal coreset and the FPT algorithm for finding an optimal assignment required to handle each of these variants, and we direct the reader to these discussions. Since these are the only two ingredients required in our problem, our framework implies an EPAS for all of these variants in metric spaces of bounded scatter dimension.
4.2.2 Vanilla (Voronoi) Clustering
@@
First, let us suppose that there exists a point $p\in W$ such that ${\sf d}(p,x_{i})>\frac{6r^{\prime}_{i}}{\varepsilon}$ for all $i\in[k]$. Then, by arguments similar to that in 15.1, we can argue that ${\sf d}(p,X)<(1+\varepsilon){\sf d}(p,O)$, contradicting the assumption that $p\in W$. This shows part (i).
@@ 4.3 Additional Remarks
Jaiswal and Kumar [JK23] (also [DGI25]), extending the work of Agrawal et al. [AIS+23] gave a general additive-$\varepsilon$ approximation-preserving reduction from constrained $(k,z)$-clustering with $m$ outliers to its analogous constrained version without outliers that has a multiplicative overhead of $\left(\frac{k+m}{\varepsilon}\right)^{O(m)}\cdot n^{O(1)}$. By plugging in our EPASes for constrained clustering into this result, we also obtain EPASes for the corresponding outlier versions.
@@
For most known metric spaces with bounded scatter dimension, the recent work of Huang et al. [HLL+25] gives coresets of size $\text{poly}(k,\varepsilon)$ for a large class of constrained $(k,z)$-clustering problems, including capacitated clustering, $(\alpha,\beta)$-fair clustering, fault-tolerant clustering, and more. To put this work into a proper context w.r.t. our EPASes, two remarks are in order.
Firstly, we note that such coresets can be used to obtain EPASes for the corresponding problems [^19] Although the following argument is very simple and should be folklore, we are unable to find a reference that explicitly describes how coresets of size independent of $\log n$ may be used to obtain EPASes for constrained $(k,z)$-clustering. Therefore, we provide the argument for the sake of completeness in order to perform a fair comparison with our results.. Let $(Y,w)$ be the coreset of size $\text{poly}(k,\varepsilon)$. We start, as usual, with the color coding step, incurring a $O(2^{O(k)}|\mathcal{I}|^{O(1)})$ factor. Then, for each $i\in[k]$, we guess the cost of the $i$th optimal cluster, up to a factor of $(1+\varepsilon)$; call this value $\tilde{C}_{i}$. Next, for each $p\in Y$ and each $i\in[k]$, we guess (up to a factor of $(1+\varepsilon)$) the portion of weight of $p$ assigned to the $i$th cluster, call this $f_{i}(p)$. Note that the number of guesses is bounded by $(\frac{\log n}{\varepsilon})^{k\cdot|Y|}$, which is FPT in $k$ and $\varepsilon$. Fix one such guess. Then, for each cluster $i\in[k]$, let $\hat{F}_{i}$ be the subset of centers $c$ of color $i$ such that $\sum_{p\in Y}{\sf d}(p,c)\cdot f_{i}(p)\approx_{1+\varepsilon}\tilde{C}_{i}$. Then, we select a “best center” from $\hat{F}_{i}$ depending upon the underling constraint (for example, for capacitated setting, we select the largest capacity center in $\hat{F}_{i}$). This yields us a candidate center set $X$. Finally, compute an optimal assignment using the assignment algorithm in time FPT in $k$ and $\varepsilon$. Finally, we return the minimum-cost solution found over all guesses. Although this strategy also yields an EPAS for many of the constrained problems in metric spaces of bounded scatter dimension, we note that the running time is $(k|Y|/\varepsilon)^{O(k|Y|)}\cdot n^{O(1)}$, and the size of coresets is $\Omega_{\varepsilon}(k^{2})$. Therefore, the dependence on $k$ of the running time is $2^{\Omega_{\varepsilon}(k^{3}\log k)}\cdot n^{O(1)}$, which is much worse compared to our running time of $2^{O_{\varepsilon}(k\log k)}\cdot n^{O(1)}$. Secondly, although the result of [HLL+25] gives coresets of size independent of $\log n$ in most known metric spaces of bounded scatter dimension, it is not known to yield such coresets for all metric spaces of bounded scatter dimension. Thus, this result does not answer the open question of [ABB+23] in its entirety.